# ESIQA: Perceptual Quality Assessment of Vision-Pro-based Egocentric Spatial Images

### v1

@@ 3.2 Performance Analysis
Table 1 shows the performance of the aforementioned 22 state-of-the-art models. It can be observed that most hand-crafted models show poor performance in evaluating the perceptual quality of egocentric spatial images in 2D viewing mode, and even perform worse for the evaluation in 3D-window and 3D-immersive modes. Compared to the hand-crafted models, deep learning-based models generally exhibit better performance across all three modes. Among all the IQA models, MAINIQA [33] achieves the best performance across three modes, demonstrating excellent prediction ability and strong generalization capabilities in different modes. Across all models, the performance for 2D-viewing-mode quality assessment is generally better than that for the 3D-window and 3D-immersive settings. The result stems from the fact that these metrics are initially designed for 2D IQA and fail to consider the stereoscopic and immersive effects inherent in 3D scenes, which significantly influence the perceptual quality. Additionally, most models exhibit slightly improved performance in predicting quality in 3D-window modes compared to 3D-immersive modes, which manifests that 3D-immersive mode presents greater challenges in predicting its QoE. In future work, we will utilize the left and right views and their disparities in spatial images to build a quality assessment model with better prediction performance and suitable for different viewing modes.
@@ 4 Conclusion
In this work, we establish the first egocentric spatial images database, termed ESIQAD, comprising 400 captured images and 100 generated images along with their MOSs under three viewing modes, i.e., 2D, 3D-window, and 3D-immersive modes. Based on our ESIQAD, we analyze the human visual perception of these images across various viewing modes and conduct a benchmark study by evaluating the performance of 22 state-of-the-art NR IQA models. The results from our benchmark study indicate that most current advanced IQA models fail to accurately assess the perceptual quality of spatial images, particularly under two 3D settings. It is worthwhile to explore the stereoscopic features of spatial images and effectively utilize disparity information to further improve the performance of egocentric spatial image quality assessment metrics in the future.

### v2

@@ 3.2.1 Evaluation on the ESIQAD
Moreover, It can be observed that most hand-crafted models show poor performance in evaluating the perceptual quality of egocentric spatial images in 2D display mode, and even perform worse for the evaluation in 3D-window and 3D-immersive modes. Compared to the hand-crafted models, deep learning-based models generally exhibit better performance across all three modes. For most benchmark models, the performance for 2D-display-mode quality assessment is generally better than that for the 3D-window and 3D-immersive settings. The experimental results in Table 2 show that the ESIQAnet narrows the performance gap across these three modes, indicating its ability to well capture the stereoscopic and immersive effects of 3D scenes, which are essential for perceptual quality. Additionally, most models exhibit slightly improved performance in predicting quality in 3D-window mode compared to 3D-immersive mode, which manifests that 3D-immersive mode presents greater challenges in predicting its QoE.
Fig. 11 compares the AUC performance of the ESIQAnet-B and other 11 outstanding benchmark IQA methods on the ESIQAD under 3D-window and 3D-immersive display modes, indicating that the proposed model significantly outperforms other benchmark methods on Different vs. Similar Analysis and Better vs. Worse Analysis criteria.
3.2.2 Evaluation on the subsets of ESIQAD
We split the ESIQAD into the subset of 400 captured images and the subset of 100 synthesized images, evaluating the performance of benchmarks and our model on each subset separately. This approach provides additional insights into the performance of IQA models in different egocentric scenarios. Since the small size of the subset of synthesized images can lead to overfitting and negatively affect model training, we employ data augmentation methods to expand this subset to 400 samples, matching the size of the subset of captured images. The experimental results are exhibited in Table 3, showing the superiority of our ESIQAnet over two subsets, followed by benchmark models MANIQA, HyperIQA, and TReS, suggesting that the ESIQAnet is a robust quality metric for distinct egocentric categories under different display modes.
The experimental results indicate that the prediction performance on the subset of captured images is relatively better than on the subset of synthesized images. The captured images from the Apple Vision Pro are more realistic, featuring natural lighting, texture, and details that provide the model with rich perceptual information. In contrast, synthesized images may lack some of these natural features, making quality assessment more challenging. Additionally, the captured images align more closely with the natural scene statistics (NSS) distribution of real-world images, making them easier for the model to characterize. However, synthesized images may introduce artifacts or features that deviate from this distribution in the synthesis process, leading to reduced prediction accuracy.
3.3 Ablation Study
To validate the effect of each component of our proposed method, we conduct ablation studies on the backbone and the three attention mechanisms on ESIQAD. The results of ablation experiments are illustrated in Table 5 and Table 4.
3.3.1 Attention Mechanisms
Cross Attention Block
The cross attention blocks in the model integrate features extracted from stereo image pairs of egocentric spatial images, utilizing their disparity to capture interactive features. To evaluate the effectiveness of cross attention blocks, we conducted ablation experiments by replacing the cross attention aggregation with simple averaging of features from the left and right views. As shown in Table 4, the cross attention block proves to be essential, enhancing the model’s performance when combined with the other two modules.
Transposed Attention Block
In the ESIQAnet, the transposed attention block is employed to facilitate global interactions across channels, enabling more comprehensive information exchange throughout the model. By introducing transposed attention blocks, we can observe a clear performance gain in SRCC and PLCC, highlighting the critical role of capturing inter-channel dependencies.
Multi-head Self Attention Block
The multi-head self attention block is utilized to capture spatial dependencies and complex local relationships within features. The results in Table 4 show that the multi-head self attention block improves the model’s ability to learn detailed and robust feature representations, resulting in enhanced performance.
3.3.2 Architecture Backbone
In the ESIQAnet, we employ multi-stage VSSD blocks that are pre-trained on ImageNet-1K [7] to extract content-aware spatial features from binocular views of egocentric spatial images. We compare our method with other pre-trained networks for spatial feature extraction, including ResNet-18 [22], ResNet-34 [22], ResNet-50 [22], Vision Transformer (ViT) [8], and Swin Transformer (SwinT) [30], all pre-trained on ImageNet. For a fair comparison, we freeze the parameters of these models during feature extraction, which is the same as the ESIQAnet model. The results illustrated in Table 5 suggest that the backbone of multi-stage VSSD blocks used in ESIQAnet surpasses the other visual feature extraction networks.
@@ 4 Conclusion
In this work, we conduct a comprehensive study for egocentric image quality assessment. We establish the first egocentric spatial images database, termed ESIQAD, comprising 500 egocentric spatial images along with their quality ratings under three common display modes, i.e., 2D, 3D-window, and 3D-immersive modes. Based on our ESIQAD, we analyze the human visual perception of these images across various display modes. Furthermore, we propose a mamba2-based multi-stage feature fusion model ESIQAnet to predict the perceptual quality of egocentric spatial images across three display modes. The results from our extensive experiments indicate that the ESIQAnet outperforms 22 state-of-the-art benchmark IQA models on the ESIQAD and its subsets, manifesting that our ESIQAnet is a robust quality metric for different egocentric categories under different display modes.
Applications. The ESIQAnet addresses a critical need in IQA research, as traditional 2D metrics fail to capture the unique features of egocentric spatial images, e.g., binocular disparity, immersion effects, and first-person perspectives, which are key aspects essential for AR/VR applications. The primary goal of ESIQAnet is to monitor and optimize QoE by accurately predicting the perceptual quality of egocentric spatial images. It can serve as a versatile tool across the media pipeline, functioning as a performance evaluation metric for codecs to optimize compression while preserving user experience. It can also guide content adaptation for various display devices and assist in real-time quality monitoring for live streaming, ensuring consistent image quality and enhancing QoE in immersive applications like gaming, virtual training, and telepresence, etc.
Limitations and Future Work. The database currently only contains egocentric spatial images captured using Apple Vision Pro and iPhone devices. Although these devices were chosen for their advanced imaging capabilities, the lack of diversity in capture devices limits the generalizability of the database. Our future works will extend the database by including egocentric spatial images captured with alternative devices, such as Meta Quest 3, Varjo XR-3, or other head-mounted displays, to improve its diversity and applicability across a broader range of devices. Furthermore, the potential impact of motion sickness in egocentric content is an important aspect to explore, especially for egocentric spatial videos viewed in immersive VR environments. It is important to study how motion sickness may affect the QoE and integrate this factor into the evaluation of egocentric spatial videos, providing a more comprehensive understanding of user experience in immersive media in future work.
Acknowledgements.
This work was supported in part by the National Key R&D Program of China under Grant 2021YFE0206700, in part by the National Natural Science Foundation of China under Grants 62401365, 62271312, 62225112, 62132006, and in part by the Shanghai Pujiang Program under Grant 22PJ1407400.
Human Research Statement
Our subjective experiments only require human participants to watch everyday materials under commercially available non-hazardous equipment for a limited time and provide their feedback. Consequently, this subjective experiment is safe and has no negative impact on the human participants. Moreover, this work does not contain any personally identifiable information or sensitive data. All data is anonymized and does not involve any infringement on personal privacy. The research primarily focuses on technology development, model testing, and algorithm optimization, with no direct impact on human subjects. The research adheres to all relevant ethical standards and legal regulations, ensuring that data usage and handling comply with ethical requirements. Therefore, IRB approval is not required in this subjective experiment. In addition, we have obtained the consent of human subjects in the study by signing an informed consent form before the subjective experiment. All participants were fully informed about the content of the experiments, their voluntary involvement, and the use of their feedback.